OpenStudy (anonymous):

A rocket, initially at rest on the ground, accelerates straight upward from rest with constant acceleration 39.2 . The acceleration period lasts for time 9.00 until the fuel is exhausted. After that, the rocket is in free fall. Find the maximum height reached by the rocket. Ignore air resistance and assume a constant acceleration due to gravity equal to.

OpenStudy (saifoo.khan):

u = 0 a = 39.2 t = 9 s = ? s = ut + 1/2at^2 s = (0)(9) + 1/2 (39.2)(9)^2 s = 3,175.2meters

OpenStudy (anonymous):

is that the final answer

OpenStudy (saifoo.khan):

Yes.

OpenStudy (anonymous):

i didnt get the same answer

OpenStudy (saifoo.khan):

Why>

OpenStudy (anonymous):

the answer is 1587